# A geometric approach to informed MCMC sampling

### v1

@@ 8 Discussion
In this work, we have proposed an original framework for developing informative MCMC schemes. The availability of explicit expressions for the exponential and inverse exponential maps under the novel use of square-root representation for pdfs plays a crucial role in constructing computationally efficient Riemannian manifold geometric MCMC algorithms. The Riemannian manifold MCMC algorithms available in the literature work for only continuous targets and involve heavy computational burden for evaluating the transition densities as well as for adjusting the tuning parameters. On the other hand, the proposed geometric MH algorithm works for both discrete and continuous targets, provides a simple step size tuning as in RWM based on acceptance rates criteria, and allows a flexible framework for using cheap to evaluate local and global approximations of the target to control the directions of moving the candidate density while still using the exact target density in the MH acceptance rate. Thus, the proposed method shows the utility of the explicit use of the intrinsic geometry of the space of pdfs in constructing informative MCMC schemes.
@@
The article presents various avenues for potential extensions and future methodological and theoretical works. The proposed geometric MCMC scheme is general and can be applied to sample from arbitrary discrete or continuous targets. The methodology is flexible and allows general choices for the baseline density and various local/global approximations of the target for specifying the directions of moving the baseline proposal density. For the popular Bayesian variable selection model considered here, we have developed some specific choices for the densities $f$ and $g$ and are implemented in the accompanying R package geommc. In particular, we have used RW base densities and the target pmf on a neighborhood as $g$. We have demonstrated the efficiency of the resulting geometric MH algorithm through analyses of high dimensional simulated datasets and a gigantic real dataset with close to a million markers for GWAS. One can consider the Locally Informed and Thresholded proposal of Zhou et al. (2022) and the adaptively scaled individual adaptation proposal of Griffin et al. (2021) as base densities. In particular, by allowing multiple variables to be added or deleted from the model in a single iteration, the algorithm can make large jumps in model space (Guan and Stephens, 2011; Liang et al., 2022). For $g$, one can consider various local tempered and non-tempered versions of $\psi$ (Zanella and Roberts, 2019). We anticipate future works exploring and comparing different choices of $\mathcal{G}$ mentioned in Section 3 and developing appropriate $f$ and $g$ functions for constructing efficient geometric MCMC schemes for other classes of statistical models.
@@
Algorithm 1 uses a fixed $\epsilon$; on the other hand, one can choose $\epsilon$ adaptively, on the fly. A potential future work is to build such adaptive versions of Algorithm 1. Also, it will be interesting to study the performance of the geometric MH algorithm when used in a Metropolis-within-Gibbs framework. This will be useful in different settings, for example, when the Bayesian variable selection example in Section 3 is extended to handle ordinal responses (Zheng et al., 2023). Zhou et al. (2022) prove that the mixing time of LIT-MH is independent of the number of covariates under the assumptions of Yang et al. (2016). Zhou and Chang (2023) derived mixing time bounds for random walk MH algorithms for high-dimensional statistical models with discrete parameter spaces and more recently Chang and Zhou (2024) extended these results to study Zanella’s (2020) informed MH algorithms. A future study is to undertake such mixing time analysis of our proposed geometric MH algorithm for Bayesian variable selection.
Appendices
Appendix A An alternative geometric MH algorithm
@@
$\tilde{\alpha}_{i^{\prime}}(x,y)=\mbox{min}\Big{\{}\frac{\psi(y)\phi_{i^{\prime},\epsilon}(x|y)}{\psi(x)\phi_{i^{\prime},\epsilon}(y|x)},1\Big{\}}.$ (24)
@@
Appendix B Further simulation results for the Bayesian variable selection example
@@
Table 5: Results for different samplers for the Bayesian variable selection example ($R^{2}=75\%$).
@@
Table 6: Results for different samplers for the Bayesian variable selection example ($R^{2}=60\%$).

### v2

@@ 8 Discussion
In this work, we have proposed an original framework for developing informative MCMC schemes. The availability of explicit expressions for the exponential and inverse exponential maps under the novel use of square-root representation for pdfs plays a crucial role in constructing computationally efficient Riemannian manifold geometric MCMC algorithms. The Riemannian manifold MCMC algorithms available in the literature work for only continuous targets and involve heavy computational burden for evaluating the transition densities as well as for adjusting the tuning parameters. On the other hand, the proposed geometric MH algorithm works for both discrete and continuous targets, provides a simple step size tuning as in RWM based on acceptance rates criteria, and allows a flexible framework for combining localized steps of the base kernel with cheap to evaluate local and global approximations of the target to control the directions of ‘global’ moves of the candidate density while still using the exact target density in the MH acceptance rate. Indeed, as demonstrated through examples, the concurrent global-local steps facilitate moves between the modes and simultaneous exploration of the modal regions. Thus, the proposed method shows the utility of the explicit use of the intrinsic geometry of the space of pdfs in constructing informative MCMC schemes.
@@
The article presents various avenues for potential extensions and future methodological and theoretical works. The proposed geometric MCMC scheme is general and can be applied to sample from arbitrary discrete or continuous targets. The methodology is flexible and allows general choices for the baseline density and various local/global approximations of the target for specifying the directions of moving the baseline proposal density. For the popular Bayesian variable selection model considered here, we have developed some specific choices for the densities $f$ and $g$ and are implemented in the accompanying R package geommc. In particular, we have used RW base densities and the target pmf on a neighborhood as $g$. We have demonstrated the efficiency of the resulting geometric MH algorithm through analyses of high dimensional simulated datasets and a gigantic real dataset with close to a million markers for GWAS. One can consider the Locally Informed and Thresholded proposal of Zhou et al. (2022) and the adaptively scaled individual adaptation proposal of Griffin et al. (2021) as base densities. In particular, by allowing multiple variables to be added or deleted from the model in a single iteration, the algorithm can make large jumps in model space (Guan and Stephens, 2011; Liang et al., 2022). For $g$, one can consider various local tempered and non-tempered versions of $\psi$ (Zanella and Roberts, 2019). We anticipate future works exploring and comparing different choices of $\mathcal{G}$ mentioned in Section 3 and developing appropriate $f$ and $g$ functions for constructing efficient geometric MCMC schemes for other classes of statistical models. These choices of $f$ and $g$ densities can then be implemented in the geommc package for specific applications.
@@
Algorithm 1 uses a fixed $\epsilon$; on the other hand, one can choose $\epsilon$ adaptively, on the fly. A potential future work is to build such adaptive versions of Algorithm 1. Also, it will be interesting to study the performance of the proposed geometric MH algorithm for the Bayesian variable selection example in the context of analyzing ordinal responses (Zheng et al., 2023), most likely extended in a Metropolis-within-Gibbs framework. Zhou et al. (2022) prove that the mixing time of LIT-MH is independent of the number of covariates under the assumptions of Yang et al. (2016). Zhou and Chang (2023) derived mixing time bounds for random walk MH algorithms for high-dimensional statistical models with discrete parameter spaces and more recently Chang and Zhou (2024) extended these results to study Zanella’s (2020) informed MH algorithms. A future study is to undertake such mixing time analysis of our proposed geometric MH algorithm for Bayesian variable selection.
Supplemental Materials
The supplemental materials contain an alternative formulation of the geometric MH framework, proofs of theoretical results, additional technical derivations for some posterior sampling, several plots corresponding to different examples considered in the main article, numerical results from an additional data analysis for the Bayesian logistic regression example, and further simulation studies on the Bayesian variable selection example.
Supplementary Materials
S1 An alternative geometric MH algorithm
@@
$\tilde{\alpha}_{i^{\prime}}(x,y)=\mbox{min}\Big{\{}\frac{\psi(y)\phi_{i^{\prime},\epsilon}(x|y)}{\psi(x)\phi_{i^{\prime},\epsilon}(y|x)},1\Big{\}}.$ (S1)
@@
S2 Proofs of theoretical results
Proof of Proposition 1.
We have
$+\sin\big{(}\epsilon\theta_{i,x}[\sin(\theta_{i,x})]^{-1}\big{\|}\sqrt{g_{i}(y|x)}-\sqrt{f(y|x)}\cos(\theta_{i,x})\big{\|}\big{)}\frac{\big{[}\sqrt{g_{i}(y|x)}-\sqrt{f(y|x)}\cos(\theta_{i,x})\big{]}}{\big{\|}\sqrt{g_{i}(y|x)}-\sqrt{f(y|x)}\cos(\theta_{i,x})\big{\|}}.$ (S2)
Note that
$\quad-2\big{\langle}\sqrt{f(y|x)},\sqrt{g_{i}(y|x)}\big{\rangle}^{2}$
$=1-\big{\langle}\sqrt{f(y|x)},\sqrt{g_{i}(y|x)}\big{\rangle}^{2}.$ (S3)
Using the fact that $\sin(\cos^{-1}(r))=\sqrt{1-r^{2}},$ we then have
$\epsilon\theta_{i,x}[\sin(\theta_{i,x})]^{-1}\big{\|}\sqrt{g_{i}(y|x)}-\sqrt{f(y|x)}\cos(\theta_{i,x})\big{\|}=\epsilon\theta_{i,x}.$ (S4)
Using (S4) in (S2), then we have
$\exp_{\sqrt{f}}(\epsilon\exp^{-1}_{\sqrt{f}}(\sqrt{g_{i}}))(y|x)=\cos(\epsilon\theta_{i,x})\sqrt{f(y|x)}+\sin(\epsilon\theta_{i,x})\zeta_{i}(y|x).$ (S5)
Next, squaring both sides of (S5) and using the fact that $2\cos(\epsilon\theta_{i,x})\sin(\epsilon\theta_{i,x})=\sin(2\epsilon\theta_{i,x})$ we have (1). Now
$\|\zeta_{i}(y|x)\|^{2}$ $=\langle\zeta_{i}(y|x),\zeta_{i}(y|x)\rangle$
$=1,$ (S6)
$\zeta_{i}\in T_{\sqrt{f}}\mathcal{M}$. Thus from (S2) it follows that (1) is a pdf. ∎
Proof of Theorem 1.
Case $c>1$:
$v(t,P)=\frac{1}{c}v(t,P_{1})+\frac{1-c}{c}\sigma^{2}_{t}.$
Since $v(t,P_{1})\leq v(t,Q)$, we have
$v(t,P)\leq\frac{1}{c}v(t,Q)+\frac{1-c}{c}\sigma^{2}_{t}.$
Case $c\leq 1$:
$v(t,Q_{1})=\frac{1}{c}v(t,Q)+\frac{1-c}{c}\sigma^{2}_{t}.$
Proof of Theorem 2.
Note that
$=c_{\epsilon}\psi(x)Q_{f}(x,dy)\mu(dx),$
Proof of Lemma 1.
$=\frac{\sin^{2}(\epsilon\theta)}{1-\exp(-1/4)}+r(x),$
where
$r(x)=\cos^{2}(\epsilon\theta)\exp(x-1/2)-[\sin^{2}(\epsilon\theta)\exp(-1/8)/\{1-\exp(-1/4)\}]\exp(x/2-1/4).$
It can be shown that $r(x)$ attains minimum at $c_{0}=2\log(\sin^{2}(\epsilon\theta)\exp(-1/8)/[2\cos^{2}(\epsilon\theta)\\ \{1-\exp(-1/4)\}])+1/2$. Thus,
which is positive for $\epsilon>0$. Thus, by Proposition 2, the geometric MH algorithm is uniformly ergodic in this example. ∎
S3 Extra plots for the Example 2
Figure S1 provides additional autocorrelation function plots for the Example 2. In particular, it shows that unlike the RW chains, the geometric MH algorithms with a diffuse normal density for $g$ lead to fast decaying autocorrelations.
Figure S1: Autocorrelation function plots of the geometric MH chains with $g$ as a diffuse normal density and (a) $t_{2}$ baseline density and (b) $N(x,1)$ baseline density where $x$ is the current state of the Markov chain.
S4 Extra plots for the Example 4
Figure S2 shows the marginal histograms obtained from the samples of the GMC($X_{2}$)-w-Gibbs chain.
Figure S2: Desired and attained marginals from 100,000 iterations of the GMC($X_{2}$)-w-Gibbs sampler for the six-mode target example. The histograms show that the chain successfully moves between the modes.
S5 Additional real data example for the Bayesian logistic regression example
[…]
Table S1: Multivariate ESS and ESS (minimum, median, maximum) and their time normalized values for different samplers for the German Credit data
Sampler mESS ESS mESS/sec ESS/sec MSJD
RW 1282 (785,1136,1622) 21 (13,19,27) 0.008
GMC(RW) 17563 (7775,17354,19004) 75 (33,74,81) 0.133
Ind 42315 (28485,41480,45140) 515 (347,505,550) 0.259
GMC(Ind) 16652 (9094,15602,18370) 81 (44,76,90) 0.128
Ind($f_{1}$) - - - - 0.000
GMC(Ind ($f_{1}$)) 15869 (7539,15542, 18804) 75 (36,73,89) 0.129
MALA 3689 (1236,2648,3880) 20 (7,14,21) 0.011
GMC(MALA) 18796 (8981,17700,20389) 33 (16,31,36) 0.138
MMALA 22938 (18817,22470,24204) 4 (4,4,5) 0.099
GMC(MMALA) 29239 (21238,28773,31873) 1 (1,1,1) 0.172
Table S2: First eight autocorrelations for different samplers for the German Credit data
Sampler lag1 lag2 lag3 lag4 lag5 lag6 lag7 lag8
RW 0.979 0.958 0.938 0.918 0.898 0.880 0.861 0.843
GMC(RW) 0.731 0.547 0.421 0.332 0.270 0.224 0.188 0.161
Ind 0.463 0.272 0.186 0.134 0.103 0.082 0.067 0.057
GMC(Ind) 0.742 0.567 0.445 0.359 0.297 0.247 0.207 0.177
Ind($f_{1}$) 0.995 0.991 0.988 0.985 0.981 0.978 0.976 0.972
GMC(Ind ($f_{1}$)) 0.742 0.568 0.449 0.364 0.303 0.253 0.214 0.184
MALA 0.949 0.901 0.856 0.815 0.776 0.739 0.705 0.672
GMC(MALA) 0.722 0.537 0.411 0.326 0.265 0.221 0.187 0.161
MMALA 0.634 0.420 0.287 0.202 0.142 0.098 0.068 0.054
GMC(MMALA) 0.594 0.358 0.219 0.135 0.084 0.053 0.033 0.025
S6 Detailed steps in MCMC sampling for the Bayesian spatial GLMM example
The joint posterior density of $(S,\beta,\sigma^{2},\theta,\omega)$ is
$\psi(s,\beta,\sigma^{2},\theta,\omega|z)=\bigg{[}\prod_{i=1}^{m}\pi(y_{i}|s_{i})\bigg{]}\pi(s|\beta,\sigma^{2},\theta,\omega)\pi(\beta|\sigma^{2})\pi(\sigma^{2})\pi(\theta)\pi(\omega),$ (S7)
where
$\pi(y_{i}|s_{i})=\frac{\exp\{-t_{i}e^{s_{i}}+s_{i}z_{i}\}t_{i}^{z_{i}}}{z_{i}!},$
$\pi(s|\beta,\sigma^{2},\theta,\omega)$ is the density of $N_{m}(X\beta,\Sigma)$, $\pi(\beta|\sigma^{2})$ is the density of $N_{p}(\mu_{0},\sigma^{2}\Sigma_{0})$, and $\pi(\sigma^{2}),\pi(\theta),\pi(\omega)$ are the densities of Inverse $\text{Inverse Gamma}(\alpha_{1},\gamma_{1}),\text{Inverse Gamma}(\alpha_{2},\gamma_{2}),$ and Inverse $\text{Inverse Gamma}(\alpha_{3},\gamma_{3})$, respectively.
To construct the various MH-within-Gibbs algorithms discussed in Example 6 of the main article for (S7), we now derive its conditional densities. First, the full conditional density of $S$ is
$\psi(s|\beta,\sigma^{2},\theta,\omega,z)\propto\bigg{[}\prod_{i=1}^{m}\frac{\exp\{-t_{i}e^{s_{i}}+s_{i}z_{i}\}t_{i}^{z_{i}}}{z_{i}!}\bigg{]}|\Sigma|^{-1/2}\exp\bigg{\{}-\frac{1}{2}(s-X\beta)^{\top}\Sigma^{-1}(s-X\beta)\bigg{\}}.$ (S8)
For constructing MALA and MMALA for the conditional distribution of $S$, we must derive the first and higher order derivatives of the logarithm of the density (S8). Since
$\log\psi(s|\beta,\sigma^{2},\theta,\omega,z)=\sum_{i=1}^{m}\bigg{\{}-t_{i}e^{s_{i}}+s_{i}z_{i}+z_{i}\log t_{i}\bigg{\}}-\frac{1}{2}\log|\Sigma|-\frac{1}{2}(s-X\beta)^{\top}\Sigma^{-1}(s-X\beta)+c_{1},$ (S9)
for some constant $c_{1}$, we have
$\frac{\partial\log\psi(s|\beta,\sigma^{2},\theta,\omega,z)}{\partial s}$ $=-t\cdot e^{s}+z-\Sigma^{-1}(s-X\beta),$
$\frac{\partial^{2}\log\psi(s|\beta,\sigma^{2},\theta,\omega,z)}{\partial s^{2}}$ $=-\text{diag}(t\cdot e^{s})-\Sigma^{-1}\equiv-G^{-1},\;\mbox{say},$
$\mbox{and}\;\frac{\partial\{G^{-1}\}_{(i,j)}}{\partial s_{k}}$ =$=\begin{cases}t_{k}\exp\{s_{k}\},&\text{iff}\;i=j=k,\\ 0,&\text{otherwise},\end{cases}$
where for two $d$ dimensional vectors $a=(a_{1},\dots,a_{d})^{\top}$ and $b=(b_{1},\dots,b_{d})^{\top}$, $a\cdot b=(a_{1}b_{1},\dots,a_{d}b_{d})^{\top}$, and $\text{diag}(a)$ denotes the $d\times d$ diagonal matrix with diagonal elements $a$.
From (S7) we see that conditional on $s$, the distributions of $\beta,\sigma^{2},\theta$ and $\omega$ do not depend on $z$. The conditional density of $\beta$ is
$\psi(\beta|s,\sigma^{2},\theta,\omega)\propto$ $\pi(s|\beta,\sigma^{2},\theta,\omega)\pi(\beta|\sigma^{2})$
$\propto$ $\exp\bigg{\{}-\frac{1}{2\sigma^{2}}(s-X\beta)^{\top}\Sigma_{\theta,\omega}^{-1}(s-X\beta)\bigg{\}}\exp\bigg{\{}-\frac{1}{2\sigma^{2}}(\beta-\mu_{0})^{\top}\Sigma_{0}^{-1}(\beta-\mu_{0})\bigg{\}},$
where $\Sigma=\sigma^{2}\Sigma_{\theta,\omega}$. Thus,
$\beta|s,\sigma^{2},\theta,\omega\sim N_{p}\bigg{\{}\bigg{(}X^{\top}\Sigma_{\theta,\omega}^{-1}X+\Sigma_{0}^{-1}\bigg{)}^{-1}\bigg{(}X^{\top}\Sigma_{\theta,\omega}^{-1}s+\Sigma_{0}^{-1}\mu_{0}\bigg{)},\sigma^{2}\bigg{(}X^{\top}\Sigma_{\theta,\omega}^{-1}X+\Sigma_{0}^{-1}\bigg{)}^{-1}\bigg{\}}.$ (S10)
Next,
$\psi(\sigma^{2}|s,\beta,\theta,\omega)\propto$ $\pi(s|\beta,\sigma^{2},\theta,\omega)\pi(\beta|\sigma^{2})\pi(\sigma^{2})$
$\propto$ $(\sigma^{2})^{-m/2}(\sigma^{2})^{-\alpha_{1}-1}(\sigma^{2})^{-p/2}\exp\bigg{\{}-\frac{\gamma_{1}}{\sigma^{2}}\bigg{\}}\exp\bigg{\{}-\frac{1}{2\sigma^{2}}(s-X\beta)^{\top}\Sigma_{\theta,\omega}^{-1}(s-X\beta)\bigg{\}}$
$\exp\bigg{\{}-\frac{1}{2\sigma^{2}}(\beta-\mu_{0})^{\top}\Sigma_{0}^{-1}(\beta-\mu_{0})\bigg{\}},$
that is,
$\sigma^{2}|s,\beta,\theta,\omega\sim\text{Inverse Gamma}\bigg{(}\frac{2\alpha_{1}+m+p}{2},\frac{2\gamma_{1}+(s-X\beta)^{\top}\Sigma_{\theta,\omega}^{-1}(s-X\beta)+(\beta-\mu_{0})\Sigma_{0}^{-1}(\beta-\mu_{0})}{2}\bigg{)}.$ (S11)
The conditional pdf of $\theta$ is
$\psi(\theta|s,\beta,\sigma^{2},\omega)\propto$ $\pi(s|\beta,\sigma^{2},\theta,\omega)\pi(\theta)$
$\propto$ $|\Sigma|^{-1/2}\exp\bigg{\{}-\frac{1}{2}(s-X\beta)^{\top}\Sigma^{-1}(s-X\beta)\bigg{\}}\theta^{-\alpha_{2}-1}\exp\bigg{\{}-\frac{\gamma_{2}}{\theta}\bigg{\}},$
which is not a standard density. Letting $\nu=\log\theta$, we use a RWM step for sampling from $\psi(\nu|s,\beta,\sigma^{2},\omega)$ which is given by
$\psi(\nu|s,\beta,\sigma^{2},\omega)\propto|\Sigma|^{-1/2}\exp\bigg{\{}-\frac{1}{2}(s-X\beta)^{\top}\Sigma^{-1}(s-X\beta)\bigg{\}}\exp\{-\alpha_{2}\nu\}\exp\bigg{\{}-\frac{\gamma_{2}}{\exp\{\nu\}}\bigg{\}}.$ (S12)
Finally, the conditional pdf of $\omega$ is
$\psi(\omega|s,\beta,\sigma^{2},\theta)$ $\propto\pi(s|\beta,\sigma^{2},\theta,\omega)\pi(\omega)$
$\propto|\Sigma|^{-1/2}\exp\bigg{\{}-\frac{1}{2}(s-X\beta)^{\top}\Sigma^{-1}(s-X\beta)\bigg{\}}\omega^{-\alpha_{3}-1}\exp\bigg{\{}-\frac{\gamma_{3}}{\omega}\bigg{\}}.$
Letting $\zeta=\log\omega$, we use a RWM step for sampling from $\psi(\zeta|s,\beta,\sigma^{2},\theta)$ given by
$\psi(\zeta|s,\beta,\sigma^{2},\theta)\propto|\Sigma|^{-1/2}\exp\bigg{\{}-\frac{1}{2}(s-X\beta)^{\top}\Sigma^{-1}(s-X\beta)\bigg{\}}\exp\{-\alpha_{3}\zeta\}\exp\bigg{\{}-\frac{\gamma_{3}}{\exp\{\zeta\}}\bigg{\}}.$ (S13)
Starting with some initial values, the different variables are sequentially updated in the following order using MH-within-Gibbs algorithms.
  • Step 1
    Update $s$ with one of the MH algorithms mentioned in Example 6 in the main article.
  • Step 2
    Update $\beta$ with a draw from (S10).
  • Step 3
    Update $\sigma^{2}$ with a draw from (S11).
  • Step 4
    Update $\theta(\nu)$ by drawing
    • (a)
      a sample $\nu^{\prime}\sim N(\nu,h_{\nu})$, and
    • (b)
      accepting $\nu^{\prime}$ with probability
      $\text{min}\bigg{\{}\frac{\psi(\nu^{\prime}|s,\beta,\sigma^{2},\omega)}{\psi(\nu|s,\beta,\sigma^{2},\omega)},1\bigg{\}},$
      where $\psi(\nu|s,\beta,\sigma^{2},\omega)$ is given in (S12).
  • Step 5
    Update $\omega(\zeta)$ by drawing
    • (a)
      a sample $\zeta^{\prime}\sim N(\zeta,h_{\zeta})$, and
    • (b)
      accepting $\zeta^{\prime}$ with probability
      $\text{min}\bigg{\{}\frac{\psi(\zeta^{\prime}|s,\beta,\sigma^{2},\theta)}{\psi(\zeta|s,\beta,\sigma^{2},\theta)},1\bigg{\}},$
      where $\psi(\zeta|s,\beta,\sigma^{2},\theta)$ is given in (S13).
The step-sizes $h_{\nu}$ and $h_{\zeta}$ are chosen to achieve certain empirical acceptance rates.
S7 Further simulation results for the Bayesian variable selection example
@@
Table S3: Results for different samplers for the Bayesian variable selection example ($R^{2}=75\%$).
@@
Table S4: Results for different samplers for the Bayesian variable selection example ($R^{2}=60\%$).